(a) d d A B C (b) d d А B сConsider three identical atoms (A, B, C) which are equally separated with a large interatomic distance d as shown in FIG. 2 (a) Write the total Hamiltonian H in the matrix form with A).UB), and uc) as a basis, and find the eigenvalues. b) Suppose tha
Answer #1

Ary (@) to calculate to matrix end its H Now total When He has perturbed Hametonian We write the following Matrix. HIB Hiz HiMBB E o to similarly < 481 481487 - Hot (46) H! Y8) <48/419) = -2 14 19x><a/4alve) <48/414187 <Yel4] get Similarly YR we CaHo-dI] -0 da to , Eo to Thero (5) mayo () May e and we can write 1407 = 1977 - 1927+14) 149% = 14.7-143) +142) YA (1:0) * 14,(© Operator D d 0 Q e o -d Eu qonvectores rs are YA YR Y TYAIDITA) 147-192) +143) i Esth142 1447t=t etto/h 14,7- 164/4 1927 tLitotth d E e * d Teselt 7/35 , 1. - LYA e olya ) -d EX] 3x) (tolags)? (1-004). d d ro CScanned with CamScanner